# On endosplit $p$-permutation resolutions and Broué’s conjecture for $p$-solvable groups

### v2

@@ 3. Endosplit $p$-permutation resolutions and relatively endotrivial complexes
Let $k$ be any field of characteristic $p>0$, and let $M$ be an endo-$p$-permutation $kG$-module. $M$ has an endosplit $p$-permutation resolution $C_{M}$ if and only if there exists a $V(\mathcal{F}_{G})$-endotrivial $kG$-module $E$ which is a tensor of relative syzygies and a permutation module $X$ such that $M$ is a direct summand of $E\otimes_{k}X$. If this occurs and if $C_{M}$ has no contractible direct summands, then $C_{M}$ is a direct summand of a chain complex $C_{E}\otimes_{k}X[0]$, where $C_{E}$ is a chain tensor of relative syzygies which resolves $E$.
@@
Now, suppose $M$ has an endosplit $p$-permutation resolution $C_{M}$. Then since $h:\mathcal{E}_{k}^{V(\mathcal{F}_{G})}\to CF(G,p)$ is surjective, there exists a $V(\mathcal{F}_{G})$-endotrivial complex $C_{E}$ for which $h_{C_{E}}(P)=h_{C_{M}}(P)$ for all $P\in s_{p}(G)$ for which $C_{M}(P)$ is not contractible, and $h_{C_{E}}(Q)=0$ for all $Q\in s_{p}(G)$ for which $C_{M}(Q)$ is contractible. Moreover, we may take $C_{E}$ to be a chain tensor of relative syzygies. Since $C_{E}$ is also an endosplit $p$-permutation resolution, the module which it resolves, $E$, is a module tensor of relative syzygies. Now, $C_{E}^{*}\otimes_{k}C_{M}$ is an endosplit $p$-permutation resolution for $E^{*}\otimes_{k}M$ which has partial h-marks entirely zero. It follows by inductive removal of contractible summands (similar to the proof of Proposition 3.3) that $C_{E}^{*}\otimes_{k}C_{M}\simeq(E^{*}\otimes_{k}M)[0]$. Set $X:=E^{*}\otimes_{k}M$, then $X$ is a $p$-permutation module since it is (as a chain complex) a direct summand of a chain complex of $p$-permutation modules.
@@
where $N$ is some $V(\mathcal{F}_{G})$-projective $p$-permutation $kG$-module. Extracting homology, we find
@@
where $K$ is a contractible complex. Note that $C_{E}\otimes_{k}K$ is also contractible. Combining these statements, $C_{M}$ is a direct summand of $C_{E}\otimes_{k}X[0]\oplus C_{E}\otimes_{k}K$, but since $C_{M}$ was assumed to have no direct summands, we conclude $C_{M}$ is a direct summand of $C_{E}\otimes_{k}X[0]$.
@@
Let $k$ be a finite field, and $M$ be an endo-$p$-permutation $kG$-module. Set $k^{\prime}:=\overline{k}$. The module $k^{\prime}\otimes_{k}M$ has an endosplit $p$-permutation over $k^{\prime}$ if and only if $M$ has an endosplit $p$-permutation resolution over $k$.
@@
By Theorem 3.17, it suffices to show that $M$ is a direct summand of $\tilde{E}\otimes_{k}\tilde{X}$. However, this follows by the Noether-Deuring theorem since $k^{\prime}\otimes_{k}M$ is a direct summand of $E\otimes_{k^{\prime}}X\cong k^{\prime}\otimes_{k}(\tilde{E}\otimes_{k}\tilde{X})$.
@@
Let $C$ be an endosplit $p$-permutation resolution of $kG$-modules. Then $C$ is a direct summand of the endosplit $p$-permutation resolution
@@
Dade described how all indecomposable endopermutation $kG$-modules arise from capped endopermutation $kG$-modules in a similar manner for $p$-groups. If $G$ is a $p$-group, [Dad78, Theorem 6.6] implies that if $M$ is an indecomposable endopermutation $kG$-module with vertex $Q$, then $M$ is a direct summand of $\operatorname{Ind}^{G}_{Q}\operatorname{Res}^{G}_{Q}M$. Moreover, [Dad78, Theorem 6.10] gives a way to decompose any capped endopermutation $kG$-module in terms of its cap. These facts along with the previous corollary implies the following.
@@
[Maz03] Let $G$ be a $p$-group. If $D_{k}(G)=D^{\Omega}(G)$, then every indecomposable endopermutation $kG$-module and every capped endopermutation $kG$-module has an endosplit $p$-permutation resolution.
This is a rediscovery of Mazza’s theorem, stating when $p$ is odd, every capped endopermutation $kG$-module has an endosplit $p$-permutation resolution. The question of whether, if $D_{k}(G)=D^{\Omega}(G)$, any endopermutation $kG$-module has an endosplit $p$-permutation resolution, is a bit more technical, since the decomposition of such modules is more cumbersome, see [Dad78, Proposition 6.13].
@@ 4. Galois descent of endosplit $p$-permutation resolutions
We adapt the notation conventions from [KL18]. For arbitrary extensions of commutative rings $R\subseteq R^{\prime}$, if $A$ is an $\mathcal{O}$-algebra, one can form the $R^{\prime}$-algebra $A^{\prime}:=R^{\prime}\otimes_{R}A$. Similarly, given any $A$-module $M$, one can form an $A^{\prime}$-module $R^{\prime}\otimes_{R}M$. In general, not every $A^{\prime}$-module arises in this way. Given a ring automorphism $\sigma$ of $R^{\prime}$ which restricts to the identity map on $R$ and an $A^{\prime}$-module $M^{\prime}$, we denote by ${}^{\sigma}M^{\prime}$ the $A^{\prime}$-module which is equal to $M^{\prime}$ over the subalgebra $1\otimes A$ of $A^{\prime}$, and such that $\lambda\otimes a\in A^{\prime}$ acts on $M^{\prime}$ as $\sigma^{-1}(\lambda)\otimes a$. Note that if $f:M^{\prime}\to N^{\prime}$ is an $A^{\prime}$-module homomorphism, then $f$ is also an $A^{\prime}$-module homomorphism $f:{}^{\sigma}M^{\prime}\to{}^{\sigma}N^{\prime}$. Thus, the Galois twist ()${}^{\sigma}(-)$ induces an $R$-linear (but in general not $R^{\prime}$-linear), exact autoequivalence of ${}_{A^{\prime}}\mathbf{mod}$.
@@
$-_{R}:{}_{A^{\prime}}\mathbf{mod}\to{}_{A}\mathbf{mod}\text{ and }R^{\prime}\otimes_{R}-:{}_{A}\mathbf{mod}\to{}_{A^{\prime}}\mathbf{mod},$
restriction and extension of scalars, respectively. These are both $R$-linear exact functors. Moreover, $R^{\prime}\otimes_{R}-$ is left adjoint to $-_{R}$. Moreover, these functors induce functors over the bounded chain complex categories $Ch^{b}({}_{A}\mathbf{mod})$ and $Ch^{b}({}_{A^{\prime}}\mathbf{mod})$ and the same adjunction holds.
@@
Let $(K,\mathcal{O},k)\subseteq(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be an extension of $p$-modular systems, i.e. a pair of $p$-modular systems such that $\mathcal{O}\subseteq\mathcal{O}^{\prime}$ and $J(\mathcal{O})\subseteq J(\mathcal{O}^{\prime})$. Furthermore, we assume that $k$ and $k^{\prime}$ are finite fields, and that $\mathcal{O}$ and $\mathcal{O}^{\prime}$ are absolutely unramified, i.e. $J(\mathcal{O})=p\mathcal{O}$ and $J(\mathcal{O}^{\prime})=p\mathcal{O}^{\prime}$. This is always achievable by setting $\mathcal{O}=W(k)$ and $\mathcal{O}^{\prime}=W(k^{\prime})$, the Witt vectors of $k$ and $k^{\prime}$. Set $d=[k^{\prime}:k]$. Then $\mathcal{O}^{\prime}$ is free of rank $d$ as an $\mathcal{O}$-module. Let $\sigma:k^{\prime}\to k^{\prime}$ be a generator of $\operatorname{Gal}(k^{\prime}/k)$. Then there exists a unique ring homomorphism $\sigma:\mathcal{O}^{\prime}\to\mathcal{O}^{\prime}$ lifting $\sigma:k^{\prime}\to k^{\prime}$, since $\mathcal{O}^{\prime}$ is absolutely unramified.
Moreover, the composite $(\mathcal{O}^{\prime}\otimes_{\mathcal{O}}-)_{\mathcal{O}}$ is naturally isomorphic to the functor $(-)^{\oplus d}$, see for instance [KL18, Proposition 6.3]. As a result, since the Krull-Schmidt theorem for $\mathcal{O}G$-modules and complexes holds, Galois descent is unique up to isomorphism. That is, if $M_{1},M_{2}$ are $\mathcal{O}G$-modules for which $\mathcal{O}^{\prime}\otimes_{\mathcal{O}}M_{1}\cong\mathcal{O}^{\prime}\otimes_{\mathcal{O}}M_{2}$, then $M_{1}\cong M_{2}$, and similarly for chain complexes. This also holds over the residue field $k$.
@@
[Mil25b, Theorem 6] Let $(K,\mathcal{O},k)\subseteq(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be an extension of $p$-modular systems with $\mathcal{O},\mathcal{O}^{\prime}$ absolutely unramified, and let $A$ be a finitely generated $\mathcal{O}$-free $\mathcal{O}$-algebra. Set $A^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}A$.
@@
    Suppose $C\in Ch^{b}({}_{A^{\prime}}\mathbf{mod})$ satisfies ${}^{\sigma}C\cong C$ for all $\sigma\in\operatorname{Gal}(k^{\prime}/k)$, where we regard $\sigma$ as the unique ring homomorphism of $\mathcal{O}^{\prime}$ lifting $\sigma\in\operatorname{Gal}(k^{\prime}/k)$. Then there exists a chain complex $\tilde{C}\in Ch^{b}({}_{A}\mathbf{mod})$ such that $\mathcal{O}^{\prime}\otimes_{\mathcal{O}}\tilde{C}\cong C$. Moreover, $\tilde{C}$ is unique up to isomorphism.
@@
First, assume ${}^{\sigma}M\cong M$. $C$ is an endosplit $p$-permutation resolution for $M$, and since ()${}^{\sigma}(-)$ is exact, ${}^{\sigma}C$ is an endosplit $p$-permutation resolution for ${}^{\sigma}M$. Therefore, $C^{*}\otimes_{k}{}^{\sigma}C$ is an endosplit $p$-permutation resolution for $M^{*}\otimes_{k}{}^{\sigma}M\cong M^{*}\otimes_{k}M\cong k\oplus N$ for some $V$-projective $N$.
Notice that $h_{C}(P)=h_{{}^{\sigma}C}(P)$ for all $P\in s_{p}(G)$. Therefore, $C^{*}\otimes_{k}{}^{\sigma}C$ has h-marks entirely 0 and homology ${}^{\sigma}M\otimes_{k}M^{*}\cong k\oplus N$. Therefore, $C^{*}\otimes_{k}{}^{\sigma}C\simeq(k\oplus N)[0]$, so $[C]=[{}^{\sigma}C]\in\mathcal{E}_{k}^{V}(G)$. However, both $C$ and ${}^{\sigma}C$ are indecomposable, and every equivalence class in $\mathcal{E}_{k}^{V}(G)$ has a unique indecomposable representative, thus $C\cong{}^{\sigma}C$. The reverse implication is clear since ()${}^{\sigma}(-)$ is exact. ∎
@@
Let $k^{\prime}/k$ be an extension of finite fields and $V\in{}_{k^{\prime}G}\mathbf{mod}$. Suppose $C$ is an indecomposable endosplit $p$-permutation for a $V$-endotrivial module $M\in{}_{k^{\prime}G}\mathbf{mod}$, i.e. $C$ is $V$-endotrivial. $M$ descends to $k$ if and only if $C$ does.
@@
Fix $(K,\mathcal{O},k)\subseteq(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ an extension of $p$-modular systems with $\mathcal{O},\mathcal{O}^{\prime}$ absolutely unramified (for instance, $\mathcal{O}:=W(k),\mathcal{O}^{\prime}:=W(k^{\prime})$). The reverse direction of the if and only if statement is clear. Suppose $M$ descends to a $kG$-module $\tilde{M}$. Then ${}^{\sigma}M\cong M$ for all $\sigma\in\operatorname{Gal}(k^{\prime}/k)$. By the previous lemma, ${}^{\sigma}C\cong C$ for all $\sigma\in\operatorname{Gal}(k^{\prime}/k)$ as well. Now, there exists a unique lift $\hat{C}$, a chain complex of $\mathcal{O}^{\prime}G$-modules, such that $k^{\prime}\otimes_{\mathcal{O}^{\prime}}\hat{C}\cong C$. It follows that ${}^{\sigma}\hat{C}\cong\hat{C}$ for all $\sigma\in\operatorname{Gal}(k^{\prime}/k)$ by unique lifting of $\sigma$. By Theorem 4.2, there exists a chain complex of $p$-permutation $\mathcal{O}G$-modules $\tilde{C}$ for which $\mathcal{O}^{\prime}\otimes_{\mathcal{O}}\tilde{C}\cong\hat{C}$. Then $k\otimes_{\mathcal{O}}\tilde{C}$. All together, we have
@@
If the previous theorem holds, then if $M$ descends to the $kG$-module $\tilde{M}$ and $C$ descends to a chain complex of $p$-permutation $kG$-modules $\tilde{C}$, then $\tilde{C}$ is an endosplit $p$-permutation resolution for $\tilde{M}$, since extension of scalars is exact. This result also holds over complete discrete valuation rings, even if we make no assumptions about absolute unramification, due to unique lifting of endosplit $p$-permutation resolutions.
@@
Let $C$ be an endotrivial complex of $kG$-modules with $H_{h_{C}(1)}(C)\cong k_{\omega}$ for some $\omega\in\operatorname{Hom}(G,\mathbb{F}_{p}^{\times}).$ Then there exists an endotrivial complex $\tilde{C}\in Ch^{b}({}_{\mathbb{F}_{p}G}\mathbf{triv})$ for which $k\otimes_{\mathbb{F}_{p}}\tilde{C}\cong C$. In particular,
@@
The first statement follows immediately from the previous results. This asserts that $\mathcal{T}\mathcal{E}_{k}(G)\cong\mathcal{T}\mathcal{E}_{\mathbb{F}_{p}}(G)$. The isomorphism $\mathcal{E}_{k}(G)\cong\mathcal{T}\mathcal{E}_{\mathbb{F}_{p}}(G)\times\operatorname{Hom}(G,k^{\times})$ follows from [Mil24, Remark 3.8], and the isomorphism $\mathcal{T}\mathcal{E}_{k}(G)\cong CF_{b}(G,p)$ follows from [Mil25a, Theorem 6.5]. ∎
@@
The fact that all endotrivial complexes live in $\mathbb{F}_{p}$ up to a twist by a degree one Brauer character is not that surprising, since they are determined almost entirely by the Borel-Smith functions which exist independently of scalars. Since endotrivial complexes are, in a sense, the one-sided analogue of a splendid Rickard equivalence, this may provide a heuristic reason for why one may place faith in the refined abelian defect group conjecture. Additionally, the previous corollary gives another proof of non-surjectivity in general of the Lefschetz homomorphism $\Lambda:\mathcal{E}_{k}(G)\to O(T(kG))$, which was first considered in [Mil24, Section 7].
@@
The forward direction is clear since extension of scalars is exact, therefore a descendant of $C$ is a resolution for a descendant of $M$. Suppose $M$ descends to $k$. By Theorem 3.17, $C$ is a direct summand of the tensor product $C_{E}\otimes_{k^{\prime}}X[0]$ of a chain tensor of relative syzygies $C_{E}$ and a permutation module $X$, regarded as a chain complex in degree zero. Now, observe $C_{E}$ and $X$ both in fact descend to $\mathbb{F}_{p}$ (this is clear from their constructions), therefore they both descend to any intermediate extension $k^{\prime}\supseteq k\supseteq\mathbb{F}_{p}$. This still holds after removing all contractible summands from $C_{E}\otimes_{k^{\prime}}X$, so we shall assume without loss of generality $C_{E}\otimes_{k^{\prime}}X$ has no contractible summands.
Similarly, $M$ is a direct summand of $E\otimes_{k^{\prime}}X$ which also descends to $k$, say with multiplicity $m$. Let $Y$ be the unique up to isomorphism $kG$-module for which $k^{\prime}\otimes_{k}Y\cong E\otimes_{k^{\prime}}X$ and let $C_{Y}$ be its corresponding endosplit $p$-permutation resolution, both of which are tensors of relative syzygies (up to removing contractibles). Then $k^{\prime}\otimes_{k}C_{Y}\cong C_{E}\otimes_{k^{\prime}}X[0]$. Note that $C_{Y}$ also contains no contractible direct summands. Let $\tilde{M}$ be the $kG$-module for which $k^{\prime}\otimes_{k}\tilde{M}\cong M$, then $\tilde{M}$ is a direct summand of $Y$, and also occurs with multiplicity $m$.
Let $Y=Y_{1}\oplus\cdots\oplus Y_{l}$ be a direct sum decomposition of $Y$ into indecomposables. By Proposition 3.4, this corresponds to a direct sum decomposition $C_{Y}=C_{Y_{1}}\oplus\cdots\oplus C_{Y_{l}}$ where each $C_{Y_{i}}$ is an endosplit $p$-permutation resolution for $Y_{i}$. Therefore,
$C_{E}\otimes_{k^{\prime}}X[0]\cong k^{\prime}\otimes_{k}(C_{Y_{1}}\oplus\cdots\oplus C_{Y_{l}})\cong k^{\prime}\otimes_{k}C_{Y_{1}}\oplus\cdots\oplus k^{\prime}\otimes_{k}C_{Y_{l}}.$
@@
Let $k$ be a finite field and set $k^{\prime}:=\overline{k}$. Suppose $M$ is an endo-$p$-permutation $kG$-module for which $k^{\prime}\otimes_{k}M$ has an endosplit $p$-permutation resolution $\hat{C}$ of $p$-permutation $k^{\prime}G$-modules. Then there exists an endosplit $p$-permutation resolution $C$ of $M$ such that $k^{\prime}\otimes_{k}C\cong\hat{C}$.

### v3

@@ 3. Endosplit $p$-permutation resolutions and relatively endotrivial complexes
Let $k$ be any field of characteristic $p>0$, and let $M$ be an endo-$p$-permutation kGkG-module. $M$ has an endosplit $p$-permutation resolution $C_{M}$ if and only if there exists a $V(\mathcal{F}_{G})$-endotrivial $kG$-module $E$ which is a tensor of relative syzygies and a permutation module $X$ such that $M$ is a direct summand of $E\otimes_{k}X$. If this occurs and if $C_{M}$ has no contractible direct summands, then $C_{M}$ is a direct summand of a chain complex $C_{E}\otimes_{k}X[0]$, where $C_{E}$ is a chain tensor of relative syzygies which resolves $E$.
@@
Now, suppose $M$ has an endosplit $p$-permutation resolution $C_{M}$. Then since $h\colon\mathcal{E}_{k}^{V(\mathcal{F}_{G})}\to CF(G,p)$ is surjective, there exists a $V(\mathcal{F}_{G})$-endotrivial complex $C_{E}$ for which $h_{C_{E}}(P)=h_{C_{M}}(P)$ for all $P\in s_{p}(G)$ for which $C_{M}(P)$ is not contractible, and $h_{C_{E}}(Q)=0$ for all $Q\in s_{p}(G)$ for which $C_{M}(Q)$ is contractible. Moreover, we may take $C_{E}$ to be a chain tensor of relative syzygies. Since $C_{E}$ is also an endosplit $p$-permutation resolution, the module which it resolves, $E$, is a module tensor of relative syzygies. Now, $C_{E}^{*}\otimes_{k}C_{M}$ is an endosplit $p$-permutation resolution for $E^{*}\otimes_{k}M$ which has partial $h$-marks entirely zero. It follows by inductive removal of contractible summands (similar to the proof of Proposition 3.3) that $C_{E}^{*}\otimes_{k}C_{M}\simeq(E^{*}\otimes_{k}M)[0]$. Set $X:=E^{*}\otimes_{k}M$, then $X$ is a $p$-permutation module since it is (as a chain complex) a direct summand of a chain complex of $p$-permutation modules.
@@
where $N$ is some $V(\mathcal{F}_{G})$-projective $p$-permutation kGkG-module. Extracting homology, we find
@@
where $K$ is a contractible complex. Note that $C_{E}\otimes_{k}K$ is also contractible. Combining these statements, $C_{M}$ is a direct summand of $C_{E}\otimes_{k}X[0]\oplus C_{E}\otimes_{k}K$, but since $C_{M}$ was assumed to have no contractible direct summands, we conclude $C_{M}$ is a direct summand of $C_{E}\otimes_{k}X[0]$.
@@
Let $k$ be a finite field, and $M$ be an endo-$p$-permutation kGkG-module. Set $k^{\prime}:=\overline{k}$. The module $k^{\prime}\otimes_{k}M$ has an endosplit $p$-permutation over $k^{\prime}$ if and only if $M$ has an endosplit $p$-permutation resolution over $k$.
@@
[…]
@@
Let $C$ be an endosplit $p$-permutation resolution with no contractible direct summands of a $kG$-module. Then $C$ is a direct summand of the endosplit $p$-permutation resolution
@@
Dade described how all indecomposable endopermutation $kG$-modules arise from capped endopermutation $kG$-modules in a similar manner for $p$-groups. If $G$ is a $p$-group, [D78a, Theorem 6.6] implies that if $M$ is an indecomposable endopermutation $kG$-module with vertex $Q$, then $M$ is a direct summand of $\operatorname{Ind}^{G}_{Q}\operatorname{Res}^{G}_{Q}M$. Moreover, [D78a, Theorem 6.10] gives a way to decompose any capped endopermutation $kG$-module in terms of its cap. These facts along with the previous corollary implies the following.
@@
[Ma03] Let $G$ be a $p$-group. If $D_{k}(G)=D^{\Omega}(G)$, then every indecomposable endopermutation $kG$-module and every capped endopermutation $kG$-module has an endosplit $p$-permutation resolution.
This is a rediscovery of Mazza’s theorem, stating when $p$ is odd, every capped endopermutation $kG$-module has an endosplit $p$-permutation resolution. The question of whether, if $D_{k}(G)=D^{\Omega}(G)$, any endopermutation $kG$-module has an endosplit $p$-permutation resolution, is a bit more technical, since the decomposition of such modules is more cumbersome, see [D78a, Proposition 6.13].
@@ 4. Galois descent of endosplit $p$-permutation resolutions
[…]
@@
[…]
restriction and extension of scalars, respectively. These are both $R$-linear exact functors. Moreover, $R^{\prime}\otimes_{R}$- is left adjoint to $_{R}$. Moreover, these functors induce functors over the bounded chain complex categories $Ch^{b}({}_{A}\mathbf{mod})$ and $Ch^{b}({}_{A^{\prime}}\mathbf{mod})$ and the same adjunction holds.
@@
Let $(K,\mathcal{O},k)\subseteq(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be an extension of $p$-modular systems, i.e. a pair of $p$-modular systems such that $\mathcal{O}\subseteq\mathcal{O}^{\prime}$ and $J(\mathcal{O})\subseteq J(\mathcal{O}^{\prime})$. Furthermore, we assume that $k$ and $k^{\prime}$ are finite fields, and that $\mathcal{O}$ and $\mathcal{O}^{\prime}$ are absolutely unramified, i.e. $J(\mathcal{O})=p\mathcal{O}$ and $J(\mathcal{O}^{\prime})=p\mathcal{O}^{\prime}$. This is always achievable by setting $\mathcal{O}=W(k)$ and $\mathcal{O}^{\prime}=W(k^{\prime})$, the Witt vectors of $k$ and $k^{\prime}$. Set $d=[k^{\prime}:k]$. Then $\mathcal{O}^{\prime}$ is free of rank $d$ as an $\mathcal{O}$-module. Let $\sigma\colon k^{\prime}\to k^{\prime}$ be a generator of $\operatorname{Gal}(k^{\prime}/k)$. Then there exists a unique ring homomorphism $\sigma\colon\mathcal{O}^{\prime}\to\mathcal{O}^{\prime}$ lifting $\sigma\colon k^{\prime}\to k^{\prime}$, since $\mathcal{O}^{\prime}$ is absolutely unramified, and the composite $(\mathcal{O}^{\prime}\otimes_{\mathcal{O}}-)_{\mathcal{O}}$ is naturally isomorphic to the functor $(-)^{\oplus d}$ since $\mathcal{O}^{\prime}$ is free over $\mathcal{O}$ of rank $d$. As a result, since the Krull-Schmidt theorem for $\mathcal{O}G$-modules and complexes holds, Galois descent is unique up to isomorphism. That is, if $M_{1},M_{2}$ are $\mathcal{O}G$-modules for which $\mathcal{O}^{\prime}\otimes_{\mathcal{O}}M_{1}\cong\mathcal{O}^{\prime}\otimes_{\mathcal{O}}M_{2}$, then $M_{1}\cong M_{2}$, and similarly for chain complexes. This also holds over the residue field $k$.
@@
[M24d, Theorem 6] Let $(K,\mathcal{O},k)\subseteq(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be an extension of $p$-modular systems with $\mathcal{O},\mathcal{O}^{\prime}$ absolutely unramified, and let $A$ be a finitely generated $\mathcal{O}$-free $\mathcal{O}$-algebra. Set $A^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}A$.
@@
    Suppose $C\in Ch^{b}({}_{A^{\prime}}\mathbf{mod})$ indecomposable satisfies $^{\sigma}C\cong C$ for all $\sigma\in\operatorname{Gal}(k^{\prime}/k)$, where we regard $\sigma$ as the unique ring homomorphism of $\mathcal{O}^{\prime}$ lifting $\sigma\in\operatorname{Gal}(k^{\prime}/k)$. Then there exists an indecomposable chain complex $\tilde{C}\in Ch^{b}({}_{A}\mathbf{mod})$ such that $\mathcal{O}^{\prime}\otimes_{\mathcal{O}}\tilde{C}\cong C$. Moreover, $\tilde{C}$ is unique up to isomorphism.
@@
[…]
@@
Let $k^{\prime}/k$ be an extension of finite fields and $V\in{}_{k^{\prime}G}\mathbf{mod}$. Suppose $C$ is an indecomposable endosplit $p$-permutation for a $V$-endotrivial module $M\in{}_{k^{\prime}G}\mathbf{mod}$, i.e. $C$ is $V$-endotrivial. Then $M$ descends to $k$ if and only if $C$ does.
@@
Fix $(K,\mathcal{O},k)\subseteq(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ an extension of $p$-modular systems with $\mathcal{O},\mathcal{O}^{\prime}$ absolutely unramified (for instance, $\mathcal{O}:=W(k),\mathcal{O}^{\prime}:=W(k^{\prime})$). The reverse direction of the if and only if statement is clear. Suppose $M$ descends to a $kG$-module $\tilde{M}$. Then $^{\sigma}M\cong M$ for all $\sigma\in\operatorname{Gal}(k^{\prime}/k)$. By the previous lemma, $^{\sigma}C\cong C$ for all $\sigma\in\operatorname{Gal}(k^{\prime}/k)$ as well. Now, there exists a unique lift $\hat{C}$, a chain complex of $\mathcal{O}^{\prime}G$-modules, such that $k^{\prime}\otimes_{\mathcal{O}^{\prime}}\hat{C}\cong C$. It follows that $^{}^{\sigma}\hat{C}\cong\hat{C}$ for all $\sigma\in\operatorname{Gal}(k^{\prime}/k)$ by unique lifting of $\sigma$. By Theorem 4.2, there exists a chain complex of $p$-permutation $\mathcal{O}G$-modules $\tilde{C}$ for which $\mathcal{O}^{\prime}\otimes_{\mathcal{O}}\tilde{C}\cong\hat{C}$. Then $k\otimes_{\mathcal{O}}\tilde{C}$. All together, we have
@@
If the previous theorem holds, then if $M$ descends to the $kG$-module $\tilde{M}$ and $C$ descends to a chain complex of $p$-permutation kGkG-modules $\tilde{C}$, then $\tilde{C}$ is an endosplit $p$-permutation resolution for $\tilde{M}$, since extension of scalars is exact. This result also holds over complete discrete valuation rings, even if we make no assumptions about absolute unramification, due to unique lifting of endosplit $p$-permutation resolutions.
@@
Let $C$ be an indecomposable endotrivial complex of $kG$-modules with $H_{h_{C}(1)}(C)\cong k_{\omega}$ for some $\omega\in\operatorname{Hom}(G,\mathbb{F}_{p}^{\times}).$ Then there exists an endotrivial complex $\tilde{C}\in Ch^{b}({}_{\mathbb{F}_{p}G}\mathbf{triv})$ for which $k\otimes_{\mathbb{F}_{p}}\tilde{C}\cong C$. In particular,
@@
The first statement follows immediately from the previous results. This asserts that $\mathcal{T}\mathcal{E}_{k}(G)\cong\mathcal{T}\mathcal{E}_{\mathbb{F}_{p}}(G)$. The isomorphism $\mathcal{E}_{k}(G)\cong\mathcal{T}\mathcal{E}_{\mathbb{F}_{p}}(G)\times\operatorname{Hom}(G,k^{\times})$ follows from [M24a, Remark 3.8], and the isomorphism $\mathcal{T}\mathcal{E}_{k}(G)\cong CF_{b}(G,p)$ follows from [M24c, Corollary 6.4]. ∎
@@
The fact that all endotrivial complexes live in $\mathbb{F}_{p}$ up to a twist by a degree one Brauer character is not that surprising, since they are determined almost entirely by the Borel-Smith functions which exist independently of scalars. Since endotrivial complexes are, in a sense, the one-sided analogue of a splendid Rickard equivalence, this may provide a heuristic reason for why one may place faith in the refined abelian defect group conjecture. Additionally, the previous corollary gives another proof of non-surjectivity in general of the Lefschetz homomorphism $\Lambda\colon\mathcal{E}_{k}(G)\to O(T(kG))$, which was first considered in [M24a, Section 7].
@@
The forward direction is clear since extension of scalars is exact, therefore a descendant of $C$ is a resolution for a descendant of $M$. Suppose $M$ descends to $k$. By Theorem 3.17, $C$ is a direct summand of the tensor product $C_{E}\otimes_{k^{\prime}}X[0]$ of a chain tensor of relative syzygies $C_{E}$ and a permutation module $X$, regarded as a chain complex in degree zero. Now, observe $C_{E}$ and $X$ both in fact descend to $\mathbb{F}_{p}$ (this is clear from their constructions), therefore they both descend to any intermediate extension $k^{\prime}\supseteq k\supseteq\mathbb{F}_{p}$. Hence the same holds for $C_{E}\otimes_{k^{\prime}}X[0]$.
Similarly, $M$ is a direct summand of $E\otimes_{k^{\prime}}X$ which also descends to $k$, say with multiplicity $m$. Let $Y$ be the unique up to isomorphism $kG$-module for which $k^{\prime}\otimes_{k}Y\cong E\otimes_{k^{\prime}}X$ and let $C_{Y}$ be its corresponding endosplit $p$-permutation resolution containing no contractible direct summands. Both complexes are tensors of relative syzygies, and we have $k^{\prime}\otimes_{k}C_{Y}\simeq C_{E}\otimes_{k^{\prime}}X[0]$.
Let $\tilde{M}$ be the $kG$-module for which $k^{\prime}\otimes_{k}\tilde{M}\cong M$, then $\tilde{M}$ is a direct summand of $Y$, and also occurs with multiplicity $m$.
Let $Y=Y_{1}\oplus\cdots\oplus Y_{l}$ be a direct sum decomposition of $Y$ into indecomposables. By Proposition 3.4, this corresponds to a direct sum decomposition $C_{Y}=C_{Y_{1}}\oplus\cdots\oplus C_{Y_{l}}$ where each $C_{Y_{i}}$ is an endosplit $p$-permutation resolution for $Y_{i}$ with no contractible summands. Therefore,
$C_{E}\otimes_{k^{\prime}}X[0]\simeq k^{\prime}\otimes_{k}(C_{Y_{1}}\oplus\cdots\oplus C_{Y_{l}})\cong k^{\prime}\otimes_{k}C_{Y_{1}}\oplus\cdots\oplus k^{\prime}\otimes_{k}C_{Y_{l}}.$
@@
Let $k$ be a finite field and set $k^{\prime}:=\overline{k}$. Suppose $M$ is an endo-$p$-permutation kGkG-module for which $k^{\prime}\otimes_{k}M$ has an endosplit $p$-permutation resolution $\hat{C}$ of $p$-permutation $k^{\prime}G$-modules. Then there exists an endosplit $p$-permutation resolution $C$ of $M$ such that $k^{\prime}\otimes_{k}C\cong\hat{C}$.